OpenStudy (anonymous):

Write the function as a piecewise function. f(x) = 4|x - 3| A) B) C) D)

OpenStudy (zarkon):

\[f(x)=\left\{\begin{matrix}4(x - 3)& \text{if} & x\ge 3 \\ -4(x - 3)& \text{if} &x<3\end{matrix}\right.\]